OpenStudy (anonymous):

Please explain to me how to simplify the expression below, and what is the simplified expression? (3x^3y^4)^2

OpenStudy (anonymous):

I am not 100% sure because that was Pre-Algebra I learned that (i am in algebra 2) but i believe there is some square rooting in there, i just don't remember where when and why. Maybe that will jog your memory ;)

OpenStudy (anonymous):

With (3x^3y^4)^2 you just square each term : 3^2 = 9 (x^3)^2 = x^3 x^3 = x^6 (y^4)^2 = y^4 y^4 = y^8 and now you gather together all the terms : (3x^3y^4)^2 = 9 x^6 y^8 Does this answer look ok?